OpenStudy (alphina):

will fan and medal ! What is the value of z for the equation fraction 1 over 3z = −fraction 5 over 6 + fraction 1 over 6z? 5 2 −2 −5

OpenStudy (alphina):

@Mehek14

OpenStudy (alphina):

if you get please explain don't just give the answer

OpenStudy (alphina):

@legomyego180

OpenStudy (legomyego180):

Could you rewrite this using the equation tool, or post a picture? I cant get a grasp of what the problem looks like.

OpenStudy (alphina):

ok

OpenStudy (alphina):

@legomyego180

OpenStudy (alphina):

better?

OpenStudy (legomyego180):

Yea, \[\frac{ 1 }{ 3 }z=-\frac{ 5 }{ 6 }+\frac{ 1 }{ 6 }z\] Lets rewrite this as: \[\frac{ 1 }{ 3 }z=-\frac{ 1 }{ 6 }z-\frac{ 5 }{ 6 }\]

OpenStudy (legomyego180):

one sec

OpenStudy (heyyouitsamonkey):

It's me alphina one sec

OpenStudy (heyyouitsamonkey):

ok

OpenStudy (legomyego180):

I've got the answer, its just a longer process

OpenStudy (heyyouitsamonkey):

ok

OpenStudy (legomyego180):

So first off you want to get ride of the one third on the left. So multiply everything on the left by 3. This gives you: \[z=\frac{ 1 }{ 2 }z-\frac{ 15 }{ 6 }\] Next you want to get everything with a Z on the same side. \[\frac{ 1 }{ 2 }z-z=\frac{ 15 }{ 6 }\] Now you need to factor out your z's on the left hand side. \[z(\frac{ 1 }{ 2 }-1)=\frac{ 15 }{ 6 }\] \[z(-\frac{ 1 }{ 2 })=\frac{ 15 }{ 6 }\] Now divide the right side by (-1/2) aka multiply it by -2 which gives your -30/6 which simplifies to -5

OpenStudy (alphina):

thanks :D